OpenStudy (anonymous):

The sum of two integers is -61. The difference of two integers is -143. What are the two integers?

OpenStudy (chriss):

x+y=-61 x-y=-143 If you add these together, you will eliminate y and have a solution for x. You can then put the solution for x back into one of the original equations and solve for y.

OpenStudy (lgbasallote):

x = first y= second x+y = -61 x-y = -143 2x = -204 x = -102 y = 39

OpenStudy (anonymous):

-102 and 41

OpenStudy (lgbasallote):

oh yeah41..bad mathing lol

OpenStudy (chriss):

Luis' method is another good way to do it. The first method used was elimination. Luis's method is substitution. Both work well. I happened to choose elimination in this case because of the ease of elimination the y variable because they were equal opposites as written.

OpenStudy (anonymous):

Thanks to everyone! You're very much appreciated :)

OpenStudy (chriss):

You're welcome :)